The class average on a statistics test is 65 marks with a standard deviation of 12 marks. What minimum mark must a student obtain to be in the top 8% of the class?
Answer by rothauserc(4718)

1 - 0.08 = 0.92
:
z-score associated with probability 0.92 is 1.40
:
(X - 65)/12 = 1.40
:
X - 65 = 12 * 1.40 = 16.8
:
X = 16.8 + 65 = 81.8 is approximately 82
:
Therefore, the minimum mark for a student to obtain to be in the top 8% of the class is 82
: